OpenStudy (anonymous):

16. Simplify the following: -2 – (7 + 10) 8 – 11 – 10 -2 – 7 + 3 17. Simplify the following: 4x – 2 – 3x 4x – (2 – 3x) 18. Evaluate the following: (-10 ÷ -2)(-2) 10 – 3 • 7 – 4(-2) + 3 19. Identify each as either prime or composite. 89 147 159 1 • 33 20. What is the prime factorization of the following: 250 1,251 99 21. Find the LCM for the following: 2 and 3 3, 4, and 5 24 and 36 22. List all the factors of each of the following: 24 77 50 64

OpenStudy (anonymous):

Too many questions. What do you actually want answered and what is it you're confused about?

OpenStudy (anonymous):

HI

OpenStudy (anonymous):

I CAN HELP WITH ONE OR TWO ANSWERS

OpenStudy (chris):

hey hunterprincess, you may have better luck if you break this up into multiple questions. I wouldn't change this one, but in general

OpenStudy (sandra):

I'll do the first one, but I agree, too many questions: 16. Simplify the following: -2 – (7 + 10) = -2 -7 - 10 = -19 8 – 11 – 10 = -13 -2 – 7 + 3 = -6

OpenStudy (anonymous):

ITS SO EASY WHY WONT U DO IT UR SELF

OpenStudy (shadowfiend):

Yeah, a bit of advice I presented earlier -- ask one question at a time, and wait for someone to help you. Then, try the other problems by yourself, and ask for help when you run into a wall again. You will definitely learn better and have less trouble this way :)

OpenStudy (shadowfiend):

sassy, please be respectful of others who are having trouble.

OpenStudy (anonymous):

HUH I SAID IM A FITH GRADER AND I CAN SOLVE IT

OpenStudy (anonymous):

Sassy doesn't help except to spam her tutoring site.

OpenStudy (anonymous):

I don't mind helping, but, I don't know what the problem is. Is it with distributing negatives, or order of operations? Do you not know what a prime number is, or is the problem not understanding prime factorization? Do you not understand the concept of multiples or is it LCM itself you are having issues with. I like helping as a teacher in training, but context is helpful.

OpenStudy (anonymous):

Prime factorization is where you take a number and break it up into prime numbers that, when multiplied together, give the original number. For example, if I wanted the prime factorization of 192, the first thing I would do is divide by 2. That would give 96. Then I'd divide by 2 again, giving 48. Again, divide 48 by 2 to give 24. Again for 12, and again for 6, and one last time for 3. Since 2 and 3 are both prime numbers, you're done. Since you divided the original number by 2 6 times and ended up with 3, the prime factorization would be 2^6*3 or 2*2*2*2*2*2*3.

OpenStudy (anonymous):

When you get a prime factorization, it's always a good idea to double check it by evaluating it and making sure you end up with the original.